OpenStudy (iwanttogotostanford):

PLZ HELP: Is a graph properly situated if the x-axis is given in numerical quantities? Yes or No/? True False

OpenStudy (iwanttogotostanford):

@imqwerty

imqwerty (imqwerty):

yeah it represents displacement you can check it by this fact- area under the curve- \(velocity \times time \) write their units- \(\large \frac{meter}{seconds} \times seconds= meters\) and meters is the unit of displacement here we won't say that it is distance because velocity is a vector

OpenStudy (gabbyalicorn):

True. The integral of the velocity vs time graph is the position vs time graph. The derivative of the position function of time is the velocity function.